Top at the SSC and LHC

Pekka K. Sinervo Department of Physics University of Toronto Toronto, CANADA M5S 1A7

November 6, 1992

Abstract

The experimental prospects for top quark studies at the Superconducting Supercollider (SSC) and the Large Hadron Collider (LHC) are reviewed. Estimates are made of the precision by which the top quark mass and branching ratios can be measured at these high energy hadron colliders, and the potential signatures for a charged Higgs boson lighter than the top quark are discussed.

References and Bibliography

The following articles, and the references contained therein, discuss many of the topics relevant to top quark physics at the SSC and LHC.

1. Compact Muon Solenoid Letter of Intent, LHCC 92-3 (October 1992).

2. SDC Technical Design Report, SDC-92-201 (April 1992).

3. Discovery limit of the charged Higgs boson via top quark decay at future hadron colliders, D. P. Roy, Phys. Lett. B283 (1992), 403.

4. W boson-top quark background to heavy Higgs boson production, G. A. Ladinsky and C.-P. Yuan, Phys. Rev. D43 (1991), 789.

5. Large Hadron Collider Workshop Proceedings, Vol II, ed. G. Jarlskog and D.Rein, ECFA 90-133 (December 1990).

6. The Total Cross Section for the Production of Heavy Quarks in Hadronic Collisions, P. Nason, S. Dawson and R. K. Ellis, Nucl. Phys. B303 (1988), 607.